# Revisit What You See: Disclose Language Prior in Vision Tokens for Efficient Guided Decoding of LVLMs

### v1

@@ 4.3 Additional analyses
To investigate the impact of each design choice, we conduct an ablation study on Qwen2.5-VL, evaluated under the CHAIR benchmark. Table 4 summarizes the results, varying three factors: (1) w/ or w/o vocabulary subset constraint, (2) layer selection strategy (all layers vs. last layer), and (3) threshold $\alpha$ for vocabulary selection. First, it is observed that selecting vision tokens from all layers achieves better performance than relying solely on the last layer on Qwen2.5-VL, suggesting that aggregating multi-layer information can be beneficial for visual grounding. We also observe that increasing $\alpha$ results in a smaller and sparser vocabulary subset, which improves the Recall score by promoting the generation of more plausible ground-truth objects, as observed with $\alpha=1\mathrm{e}{-4}$. However, as enlarging subset can include noisy tokens and result in the reduced precision, which indicates the importance of proper selection.
Table 4: Ablation study. We evaluate the impact of vocabulary subset constraint, layer selection, and threshold $\alpha$ variation on hallucination mitigation under the CHAIR benchmark. Main table settings are in gray.
Subset Layer $\boldsymbol{\alpha}$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
w/o all $1\mathrm{e}{-5}$ 0.2 7.14 0.83
w/ all $1\mathrm{e}{-4}$ 31.8 7.93 72.17
$1\mathrm{e}{-5}$ 31.6 7.56 70.83
$1\mathrm{e}{-6}$ 32.4 9.38 70.44
last $1\mathrm{e}{-4}$ 34.6 7.89 71.27
$1\mathrm{e}{-5}$ 33.8 7.92 70.31
$1\mathrm{e}{-6}$ 33.2 9.07 70.83
As shown in Table 4, when vision token projection is performed over the full vocabulary without constraining the subset (w/o subset), the CHAIRS score is abnormally low (0.2) and the CHAIRI score also appears improved compared to the best setting. However, the Recall score drops sharply to 0.83, indicating that the model fails to generate not only hallucinated objects but also a significant portion of ground-truth objects. This highlights the necessity of constraining the candidate set to context-relevant tokens. Examples of the generation failures are provided in Appendix C.2.
Applicability to MoE LVLM.
To further demonstrate the model-agnostic applicability of our decoding strategy, we evaluate ReVisiT on the recent DeepSeekVL2 model [42], a 27B-parameter LVLM with a Mixture-of-Experts (MoE) decoder backbone based on DeepSeek-MoE. Despite its large total parameter count, only 4.1B parameters are activated per inference due to expert sparsity. This architecture poses a distinct challenge compared to dense decoder models such as LLaVA or Qwen. During implementation, we observed frequent degeneration in the form of repetition even under vanilla greedy decoding. To ensure fair comparison, we apply a repetition penalty of $1.1$ across all methods for this experiment, following practices from prior work on decoding calibration.
Table 5: DeepSeekVL2 results on CHAIR benchmark. ReVisiT effectively reduces hallucinations while maintaining comparable recall.
Method CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
Greedy 20.40 7.04 59.51
ReVisiT 14.20 2.10 57.49
As shown in Table 5, ReVisiT consistently reduces hallucinations in the CHAIR benchmark: CHAIRS drops from 20.4 to 14.2, and CHAIRI from 7.04 to 2.1. These results reaffirm the effectiveness of ReVisiT in mitigating hallucinated object mentions across diverse LVLM architectures, including sparsely activated expert-based models.
Qualitative analysis.
To illustrate the qualitative improvements, we present examples from the LLaVA-Bench-In-The-Wild dataset, which consists of diverse real-world images requiring complex reasoning beyond conventional benchmarks. As shown in Figure 4, the input image is a cartoon illustration requiring contextual and conceptual understanding. While baseline methods such as vanilla greedy decoding fail to capture the main theme and generate overly generic or literal descriptions, or even getting stuck in repetitive phrasing. In contrast, ReVisiT successfully identifies the underlying concept: about neural network layers. More qualitative samples are provided in Appendix C.3
5 Related works
@@ 5.1 Pre-alignment
Recent advances in LVLMs have explored a range of pre-alignment strategies, in which visual and textual modalities are aligned during training to enhance multimodal understanding. Most existing approaches adopt a combination of instruction tuning and architectural refinements to achieve improved visual grounding. For instance, Qwen-VL [10] and InternVL [11] retain decoder-only LLM backbones while incorporating high-resolution vision encoders and cross-modal projection modules. These models leverage large-scale curated image-text data and modality-specific tuning objectives, leading to strong general-purpose performance across a variety of vision-language tasks. Building upon these designs, Qwen2.5-VL [13] introduces a Dynamic High-Resolution (DHR) processing pipeline and rotary position encodings tailored for vision-language alignment, further enhancing the model’s ability to process diverse image resolutions efficiently. NVLM [30] explores a broader architectural design space by implementing decoder-only, cross-attention, and hybrid variants, along with a DHR input pipeline and tile tagging mechanisms. On the other hand, there are another line of works designed to mitigate the architectural challenges within the pre-alignment space. Ferret [31] focuses on fine-grained spatial referring and grounding by introducing hybrid region representations and a spatial-aware visual sampler, enabling the model to handle points, boxes, and free-form regions. In a different direction, Ovis [32] addresses the structural mismatch between continuous visual embeddings and discrete textual embeddings. Specifically, Ovis introduces a learnable visual embedding table and a probabilistic vision tokenization scheme that mimics the structure of language token embeddings, thereby aligning the two modalities at the representation level. While these pre-alignment strategies have proven effective, they typically require extensive retraining and are not designed to enable inference-time intervention based on visual context.
5.2 Post-alignment
@@
Post-alignment refers to strategies that operate on the generated output sequence, aiming to correct hallucinations or factual inconsistencies after the generation process is complete. Unlike pre- or intra-alignment methods that intervene during training or decoding, post-alignment introduces no changes to the model’s parameters or inference-time logic. A key distinction among post-alignment methods lies in how the correction is applied. Some approaches train an auxiliary revision module, while others rely on external vision models to validate and refine the generated content. For example, LURE [21] learns a dedicated revisor that takes the original image and the LVLM-generated caption as input, and outputs a hallucination-corrected caption. The training signal is derived from statistical indicators such as co-occurrence bias and positional uncertainty observed in model outputs. In contrast, Woodpecker [22] does not require finetuning, but instead performs post-hoc correction through a multi-stage pipeline. Starting from the generated text, it extracts visual concepts, formulates corresponding questions, and verifies them using a pretrained VQA model. Each stage of this process relies on existing pretrained components including question generation, visual verification, and text editing modules to systematically identify and correct hallucinated content. While post-alignment strategies offer a practical way to improve output consistency without altering the base model, they typically rely on additional training or auxiliary models.
5.3 Intra-alignment

### v2

@@ 4.3 Additional analyses
We ablate four factors on Qwen2.5-VL under the CHAIR benchmark (Figure 7): (1) vision-token selection criterion (min-/random/max-JSD), (2) vocabulary constraint (subset vs. full), (3) layer scope (all layers vs. last layer), and (4) APC threshold $\alpha$.
Selection. Under a constrained (subset) vocabulary, max-JSD collapses (F1 =0.67=0.67) and random underperforms (F1 =17.63=17.63), while our min-JSD attains the best performance (F1 =81.16=81.16), substantiating the effectiveness of a context-aware constrained divergence for token selection (Table 10(a)).
Vocabulary constraint. Even with the same min-JSD selection, projecting over the full vocabulary fails (F1 =1.52=1.52), exhibiting a degenerate behavior where vision-token projections are referenced off-target. This confirms that constraining the candidate set to context-relevant tokens is crucial (Table 10(a)); qualitative failures are provided in Appendix C.2.
Layer scope & threshold. Across $\alpha$, the all-layers variant peaks at $\alpha=10^{-5}$ (F1 =81.16=81.16), while the last-layer variant peaks at $\alpha=10^{-3}$ (F1 =80.97=80.97). Overall sensitivity to $\alpha$ is modest (variations within $\approx\!1.5$ F1 around each peak), indicating that all-layers + constrained vocabulary + min-JSD is a robust setting (Table 10b).
5 Conclusion
In this paper, we revisited the role of vision tokens in LVLM decoding and provided quantitative evidence that they encode interpretable semantics that become explicit under semantically coherent vocabulary constraints. Building on these findings, we introduced ReVisiT, a training-free, model-agnostic decoding strategy that references the most relevant vision token via a context-aware constrained divergence and refines token logits accordingly. As highlighted in Fig. 1, ReVisiT consistently improves performance across six model sizes spanning three architectures, demonstrating scalability and model-agnostic effectiveness.
Appendix A Detailed Discussions
A.1 Motivation
A.1.1 Metric formulation (at-least-one recall@k)
Given any probability distribution $p_{\theta}(h,\mathcal{V})$ over the vocabulary $\mathcal{V}$,
let $\mathrm{TopK}_{k}(p_{\theta})\subset\mathcal{V}$ denote the set of its top-$k$ tokens by probability, and let $G\subseteq\mathcal{V}$ be the set of ground-truth (GT) object synonyms relevant to the evaluation unit.
We define the success indicator
$I(k)\;=\;\mathbf{1}\!\left\{\,\mathrm{TopK}_{k}(q)\;\cap\;G\;\neq\;\varnothing\,\right\},$ (7)
and report the mean of $I(k)$ over all evaluation units in the set $\mathcal{U}$:
$\widehat{\mathrm{Recall}}(k)\;=\;\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}I_{u}(k).$ (8)
A.1.2 Figure 3
CHAIR Rohrbach et al. (2018) identifies 275 hallucinated words across 500 captions. Among these, we focus on the subset where the hallucinated word and at least one ground-truth (GT) synonym are both representable as single tokens by the tokenizer, resulting in $n=190$ steps available for step-level analysis. Table 6 reports the recall rates with 95% Wilson score intervals.
Table 6: At-least-one GT recall at hallucinated steps ($n=190$) with 95% Wilson intervals.
Top-$k$ % of vocab at-least-one recall@k [95% CI]
50 0.033% 63.2% [56.0, 70.0]
100 0.066% 73.7% [66.9, 79.6]
500 0.33% 95.8% [91.6, 98.1]
A.1.3 Table 1
Setup formulation.
We follow the notation in Sec. 3: at decoding step $t$, the LVLM outputs $p_{\theta}(h_{T+t-1}^{L},\mathcal{V})$ as in Eq. equation 1. For CHAIR, let $\Omega$ denote the MSCOCO synonym set ($|\Omega|=403$). To ensure the clarity of object representation and avoid ambiguity from multi-token variants, we construct the single-token subset
$\Omega^{(1)}\;=\;\bigl\{\,w\in\Omega\;:\;\texttt{tokenize}(\texttt{"\textvisiblespace{}}w\texttt{") has length }1\,\bigr\},$ (9)
which contains 176 word types covering 62 object categories. At step $t$, we define the CHAIR-based ground-truth and hallucinated synonym subsets over $\Omega^{(1)}$ as $\mathcal{G}_{t}\subseteq\Omega^{(1)}$ and $\mathcal{H}_{t}=\Omega^{(1)}\setminus\mathcal{G}_{t}$. For step-level analyses, we further require that (i) the hallucinated word at $t$ is single-tokenizable and (ii) $\mathcal{G}_{t}\neq\varnothing$, yielding $n=190$ paired steps (selection details match Appendix A.1.2). We evaluate two conditions: w/ image (vision tokens present) and w/o image (all vision tokens removed); in the latter, steps are aligned by teacher forcing with the previously generated caption while keeping all other decoding details identical.
Statistical test and reporting.
For paired step-level comparisons, we compute the maxima
$m_{t}^{\mathrm{GT}}=\max_{w\in\mathcal{G}_{t}}p_{\theta}(w\mid h_{T+t-1}^{L},\mathcal{V}),\qquad m_{t}^{\mathrm{Hal}}=\max_{w\in\mathcal{H}_{t}}p_{\theta}(w\mid h_{T+t-1}^{L},\mathcal{V}),$
and form
$d_{t}^{\mathrm{with}}=m_{t}^{\mathrm{GT}}-m_{t}^{\mathrm{Hal}},\qquad d_{t}^{\mathrm{w/o}}=m_{t}^{\mathrm{GT,w/o}}-m_{t}^{\mathrm{Hal,w/o}},\qquad\Delta_{t}=d_{t}^{\mathrm{with}}-d_{t}^{\mathrm{w/o}}.$
We apply the Wilcoxon signed-rank test (two-sided; Pratt) to $d_{t}^{\mathrm{with}}$, $d_{t}^{\mathrm{w/o}}$, and $\Delta_{t}$ on the intersection of steps available in both conditions. We report the paired sample size $N$, the median differences, and Holm-adjusted $p$-values for $\Delta_{t}$.
Table 7: Wilcoxon signed-rank results at the step level. $\Delta_{t}=(\mathrm{GT}-\mathrm{Hal})_{\mathrm{with}}-(\mathrm{GT}-\mathrm{Hal})_{\mathrm{w/o}}$. NN is the number of paired steps. Holm-adjusted $p$-values are shown for $\Delta_{t}$.
[…]
All tokens 88,706 3.289-3.289 4.826-4.826   0.715\;\;0.715 $10^{-300}$
All tokens (in hallucinated set) 35,116 3.720-3.720 5.273-5.273   0.774\;\;0.774 $1.09{\times}10^{-116}$
Hallucinated steps 190 8.281-8.281 9.435-9.435 0.180-0.180 $0.95$
In the full-token analyses, the median $\Delta$ is positive and highly significant, indicating a systematic shift of the step-level ordering toward GT when vision tokens are present. For the hallucinated-step subset (n=190n{=}190), the median $\Delta$ is small and not significant, i.e., the reversed ordering (Hal >> GT) persists at those positions despite the image.
A.1.4 Figure 4
Setup formulation.
We follow Eq. equation 4 with the final decoder layer (j=Lj{=}L). For each of the 500 MSCOCO validation images, we project every vision token’s final-layer hidden state $h_{i}^{L}$ through the language-modeling head to obtain a distribution over the vocabulary,
$q_{i}(w)\;=\;p_{\theta}(h_{i}^{L},\mathcal{V})\quad(i=1,\dots,|\mathbf{v}|),$
yielding 173,173173{,}173 vision-token projections in total. We evaluate two settings on the same projections: (i) full vocabulary $\mathcal{V}$ and (ii) a CHAIR-derived object subset $\Omega^{(1)}$ of 176 single-token synonyms covering 62 MSCOCO categories denoted as Eq. 9
Main results with confidence intervals.
Table 8 summarizes token-level at-least-one recall@k with 95% Wilson intervals.
Table 8: At-least-one GT recall over N=173,173N{=}173{,}173 vision-token projections with 95% Wilson intervals. GT targets are restricted to the same 62 categories in both settings.
Top-$k$ Full vocabulary Constrained subset (176)
Top-1 2.03% [1.96, 2.10] 40.44% [40.21, 40.67]
Top-5 7.87% [7.74, 8.00] 69.56% [69.34, 69.78]
Top-10 12.38% [12.23, 12.54] 79.78% [79.59, 79.97]
Top-30 21.72% [21.53, 21.91] 89.24% [89.09, 89.39]
A.2 Dataset and evaluation setup
CHAIR evaluation.
CHAIR (Rohrbach et al., 2018) is a generative benchmark designed to measure object hallucinations in image captioning. A hallucinated object refers to any entity mentioned in the generated caption that is not present in the corresponding image. This metric is widely adopted to evaluate the visual grounding capability of vision-language models. We follow the standard CHAIR evaluation procedure based on the MS COCO annotations and the official implementation. [^2] https://github.com/LisaAnne/Hallucination Each image is associated with a set of ground-truth objects defined by the 80 MS COCO detection categories. During evaluation, noun phrases extracted from generated captions are matched against this predefined object list; mentions of out-of-vocabulary objects are ignored, while in-vocabulary objects absent from the ground-truth annotations are considered hallucinated.
Protocol. We use the official CHAIR code and MS COCO annotations, and evaluate on 500 images randomly sampled from the COCO validation set with a fixed seed (seed=42) for reproducibility, prompting each image with “Please describe this image in detail.”.
Metrics. We report the two standard CHAIR metrics and an additional F1:
$\begin{split}\text{CHAIR}_{S}&=\frac{\left|\{\text{hallucinated captions}\}\right|}{\left|\{\text{total captions}\}\right|},\\ \text{CHAIR}_{I}&=\frac{\left|\{\text{hallucinated object mentions}\}\right|}{\left|\{\text{total object mentions}\}\right|},\\ \text{F1}&=\frac{\sum_{\text{images}}\text{F1}_{\text{sample}}}{\left|\{\text{total captions}\}\right|}.\end{split}$
[…]
POPE evaluation.
POPE (Li et al., 2023b) is a discriminative benchmark designed to evaluate the object-level visual grounding capabilities of LVLMs. It formulates the task as a binary object presence question-answering problem, where the model is asked to determine whether a specific object is present in a given image. Evaluation is conducted using standard classification metrics derived from the confusion matrix, including accuracy, precision, and F1 score. Accuracy measures the overall proportion of correct predictions, while precision captures the fraction of predicted positive instances that are indeed correct. F1 score, defined as the harmonic mean of precision and recall, provides a balanced assessment of model performance on this binary classification task. The benchmark covers three datasets: MS COCO (Lin et al., 2014), A-OKVQA (Schwenk et al., 2022), and GQA (Hudson & Manning, 2019). For each dataset, POPE defines three types of query strategies—random, popular, and adversarial—yielding a total of nine evaluation scenarios. Binary questions are constructed in the form “Is there a <object> in the image?”, with the queried object selected according to the designated strategy. Random queries are uniformly sampled from the dataset’s object vocabulary. Popular queries focus on frequently occurring objects, while adversarial queries target objects that are semantically plausible in the given context but are absent from the image, thus probing the model’s reliance on prior co-occurrence statistics rather than visual evidence. Ground-truth labels indicating object presence are provided by the official POPE benchmark. [^3] https://github.com/AoiDragon/POPE For MS COCO, labels are derived from object detection annotations, whereas for A-OKVQA and GQA, they are obtained via SEEM-based segmentation annotations. Each query type consists of 3,000 binary QA instances, evenly split between positive and negative samples. As a result, we evaluate across a total of nine scenarios comprising 27,000 QA examples.
HallusionBench evaluation.
HallusionBench (Guan et al., 2024) is a diagnostic benchmark for image–context reasoning that disentangles Language Hallucination (over-reliance on parametric priors) and Visual Illusion (misinterpretation of visual inputs). All questions are binary (Yes/No) and organized into control pairs to test consistency across visual contexts. We focus on the Visual Dependent (VD) split, which comprises 591 Yes/No questions that require the image to answer; each figure consists of an original image and one or more human-edited variants (e.g., flipping, order reversing, masking, optical-character editing, object editing, color editing) to stress visual robustness. We use the official data and evaluation scripts. [^4] https://github.com/tianyi-lab/HallusionBench
Protocol. Following the official setup, we adopt the text-only GPT-4–assisted judging scheme. Let figures be indexed by $i\in\mathbb{I}$. For each figure $i$, let $\{I_{i,j}\}_{j\in\mathcal{J}_{i}}$ denote its image variants with $j=0$ the original and $j\geq 1$ edited versions, and let $\{q_{i,k}\}_{k\in\mathcal{K}_{i}}$ be the associated Yes/No questions (shared across variants). Given a model’s raw response $\mathcal{M}(I_{i,j},q_{i,k})$, the judge (GPT-4, temperature 0) receives the prompt with $q_{i,k}$ (question), $y_{i,j,k}$ (reference answer), $\mathcal{M}(I_{i,j},q_{i,k})$ (model response)} and returns a label in {Correct, Incorrect, Uncertain}. We binarize correctness as
[…]
Metrics. Let $\mathcal{V}=\{(i,j,k):i\in\mathbb{I},\,j\in\mathcal{J}_{i},\,k\in\mathcal{K}_{i}\}$ be all VD image–question pairs, and let $\mathbb{Q}=\{(i,k):i\in\mathbb{I},\,k\in\mathcal{K}_{i}\}$ be the set of control question pairs. We report the standard HallusionBench accuracies:
$\text{aAcc}\;=\;\frac{1}{|\mathcal{V}|}\sum_{(i,j,k)\in\mathcal{V}}b_{\mathcal{M}}(I_{i,j},q_{i,k}),$
$\text{qAcc}\;=\;\frac{1}{|\mathbb{Q}|}\sum_{(i,k)\in\mathbb{Q}}\mathbf{1}\!\Big\{\,\bigwedge_{j\in\mathcal{J}_{i}}b_{\mathcal{M}}(I_{i,j},q_{i,k})=1\,\Big\},$
$\text{fAcc}\;=\;\frac{1}{\sum_{i\in\mathbb{I}}|\mathcal{J}_{i}|}\sum_{i\in\mathbb{I}}\sum_{j\in\mathcal{J}_{i}}\mathbf{1}\!\Big\{\,\bigwedge_{k\in\mathcal{K}_{i}}b_{\mathcal{M}}(I_{i,j},q_{i,k})=1\,\Big\}.$
Here, aAcc is accuracy per question (overall), qAcc is accuracy per question pair (the same question must be correct across the figure’s image variants), and fAcc is accuracy per image (all questions assigned to a given image—original or edited—must be correct). Following the leaderboard convention, we additionally report Easy aAcc (restricted to original images, j=0j{=}0) and Hard aAcc (restricted to edited images, $j{\geq}1$).
MMMU evaluation.
MMMU (Yue et al., 2024) is a large-scale, multi-discipline VQA benchmark that evaluates college-level multimodal understanding and reasoning across six disciplines, 30 subjects, and heterogeneous image types (charts, diagrams, tables, maps, music sheets, chemical structures, etc.). We use the validation split for controlled comparisons. The validation split includes open-ended and multiple-choice (MC) questions. Since our LLaVA-1.5 configuration supports only single-image inputs, we exclude the 43 multi-image items from the 900 validation questions and report results on the remaining 857 items. Following the official setup, [^5] https://github.com/MMMU-Benchmark/MMMU we append instruction postfixes: for open-ended questions we add “Answer the question using a single word or phrase.”; for MC questions we add “Answer with the option’s letter from the given choices directly.” to adopt official parsing tool.
Metrics. We report overall Accuracy and additionally Subject Win/Tie vs. Greedy, which compares per-subject accuracy against a vanilla greedy-decoding baseline. Let $\mathcal{D}$ be the evaluated set (857 items after filtering) and $y(x)$ the ground-truth answer for item $x\in\mathcal{D}$. For a method $\mathcal{M}$,
$\text{Accuracy}\;=\;\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\mathbf{1}\{\,\mathcal{M}(x)=y(x)\,\}.$
Let $\mathcal{S}$ denote the set of subjects and $\mathcal{D}_{s}\subset\mathcal{D}$ the items belonging to subject $s$. Per-subject accuracy is $\text{Acc}_{\mathcal{M}}(s)=\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}\mathbf{1}\{\,\mathcal{M}(x)=y(x)\,\}.$ We then count
$\text{Win/Tie}\;=\;\big|\{\,s\in\mathcal{S}:\;\text{Acc}_{\mathcal{M}}(s)\geq\text{Acc}_{\text{Greedy}}(s)\,\}\big|.$
VQAv2 evaluation.
VQAv2 (Goyal et al., 2017) is a large-scale open-ended visual question answering benchmark where each question is annotated with 10 human answers. We use the official evaluation protocol and code. [^6] https://visualqa.org/evaluation.html The metric is designed to be robust to inter-annotator variability by averaging machine accuracy over all $10\choose 9$} subsets of annotators. Concretely, the per-question accuracy assigned to a model answer ans is
$\text{Acc}(\textit{ans})\;=\;\min\!\left\{\frac{\#\text{ of humans that said }\textit{ans}}{3},\,1\right\},$
and the dataset accuracy is the mean of $\text{Acc}(\textit{ans})$ over questions.
LLaVA-Bench-In-the-Wild dataset.
We additionally employ the LLaVA-Bench-In-the-Wild (Liu et al., 2023) dataset to qualitatively evaluate the visual grounding ability of LVLMs in open-ended image understanding. The dataset consists of 24 diverse images, including both real-world photographs and abstract illustrations, and 60 textual prompts designed to elicit complex visual reasoning and language generation. Following prior work Leng et al. (2024), we present representative qualitative examples to illustrate how different decoding strategies affect grounding fidelity and response quality in challenging scenarios.
A.3 Related works
@@ A.3.1 Pre-alignment
Recent advances in LVLMs have explored a range of pre-alignment strategies, in which visual and textual modalities are aligned during training to enhance multimodal understanding. Most existing approaches adopt a combination of instruction tuning and architectural refinements to achieve improved visual grounding. For instance, Qwen-VL Bai et al. (2023) and InternVL Chen et al. (2024c) retain decoder-only LLM backbones while incorporating high-resolution vision encoders and cross-modal projection modules. These models leverage large-scale curated image-text data and modality-specific tuning objectives, leading to strong general-purpose performance across a variety of vision-language tasks. Building upon these designs, Qwen2.5-VL Bai et al. (2025) introduces a Dynamic High-Resolution (DHR) processing pipeline and rotary position encodings tailored for vision-language alignment, further enhancing the model’s ability to process diverse image resolutions efficiently. NVLM Dai et al. (2024) explores a broader architectural design space by implementing decoder-only, cross-attention, and hybrid variants, along with a DHR input pipeline and tile tagging mechanisms. On the other hand, there are another line of works designed to mitigate the architectural challenges within the pre-alignment space. Ferret You et al. (2024) focuses on fine-grained spatial referring and grounding by introducing hybrid region representations and a spatial-aware visual sampler, enabling the model to handle points, boxes, and free-form regions. In a different direction, Ovis Lu et al. (2024) addresses the structural mismatch between continuous visual embeddings and discrete textual embeddings. Specifically, Ovis introduces a learnable visual embedding table and a probabilistic vision tokenization scheme that mimics the structure of language token embeddings, thereby aligning the two modalities at the representation level. While these pre-alignment strategies have proven effective, they typically require extensive retraining and are not designed to enable inference-time intervention based on visual context.
A.3.2 Post-alignment
@@
Post-alignment refers to strategies that operate on the generated output sequence, aiming to correct hallucinations or factual inconsistencies after the generation process is complete. Unlike pre- or intra-alignment methods that intervene during training or decoding, post-alignment introduces no changes to the model’s parameters or inference-time logic. A key distinction among post-alignment methods lies in how the correction is applied. Some approaches train an auxiliary revision module, while others rely on external vision models to validate and refine the generated content. For example, LURE Zhou et al. (2023) learns a dedicated revisor that takes the original image and the LVLM-generated caption as input, and outputs a hallucination-corrected caption. The training signal is derived from statistical indicators such as co-occurrence bias and positional uncertainty observed in model outputs. In contrast, Woodpecker Yin et al. (2024) does not require finetuning, but instead performs post-hoc correction through a multi-stage pipeline. Starting from the generated text, it extracts visual concepts, formulates corresponding questions, and verifies them using a pretrained VQA model. Each stage of this process relies on existing pretrained components including question generation, visual verification, and text editing modules to systematically identify and correct hallucinated content. While post-alignment strategies offer a practical way to improve output consistency without altering the base model, they typically rely on additional training or auxiliary models.
A.3.3 Intra-alignment